# Adaptive Learning Strategies for Mitotic Figure Classification in MIDOG2025 Challenge

### v1

@@ Abstract
Atypical mitotic figures (AMFs) are clinically relevant indicators of abnormal cell division, yet their reliable detection remains challenging due to morphological ambiguity and scanner variability. In this work, we investigated three variants of adapting the pathology foundation model UNI2-h for the MIDOG2025 Track 2 challenge. Starting from a LoRA-based baseline, we found that visual prompt tuning (VPT) substantially improved generalization, and that further integrating test-time augmentation (TTA) with Vahadane and Macenko stain normalization provided the best robustness. Our final submission achieved a balanced accuracy of 0.8837 and an ROC-AUC of 0.9513 on the preliminary leaderboard, ranking within the top 10 teams. These results demonstrate that prompt-based adaptation combined with stain-normalization TTA offers an effective strategy for atypical mitosis classification under diverse imaging conditions.
@@ 0.3 Implementation Details
During inference, we adopt a test-time augmentation (TTA) strategy to improve the robustness of predictions. Each test patch is evaluated under multiple transformations, including horizontal and vertical flips as well as $90^{\circ}$ rotations. In addition to geometric augmentations, we apply two complementary stain normalization pipelines, namely Vahadane [10] and Macenko [11]. For each input, predictions across all augmented versions are averaged to obtain the final probability. This ensemble-like strategy reduces prediction variance and improves generalization across scanners.
@@ Results
VPT + UNI2-h 0.8711 0.9014 0.8408 0.9483
VPT + UNI2-h + TTA 0.8837 0.9577 0.8097 0.9513
We compared three variants of our framework on the MIDOG2025 Track 2 dataset. As shown in Table 1, LoRA + UNI2-h baseline achieved a balanced accuracy of 0.8305 and an ROC-AUC of 0.9364. Replacing LoRA with Visual Prompt Tuning (VPT) improved the results, yielding a balanced accuracy of 0.8711 and an ROC-AUC of 0.9483. Finally, adding test-time augmentation (TTA) with Vahadane and Macenko stain normalization further increased robustness, achieving a balanced accuracy of 0.8837 and the highest ROC-AUC of 0.9513. Based on these results, we selected the TTA-enhanced VPT + UNI2-h variant as our final submission.
@@ Conclusion
In this study, we investigated three variants of adapting the UNI2-h foundation model for atypical mitosis classification in the MIDOG2025 Track 2 challenge. The LoRA-based approach provided a competitive baseline, while Visual Prompt Tuning (VPT) substantially improved generalization. Further integrating test-time augmentation (TTA) with Vahadane and Macenko stain normalization yielded the best results, achieving a balanced accuracy of 0.8837 and an ROC-AUC of 0.9513 on the preliminary test set. These findings demonstrate that prompt-based adaptation combined with stain-normalization TTA is an effective strategy for robust mitosis classification under diverse imaging conditions.

### v2

@@ Abstract
Atypical mitotic figures (AMFs) are clinically relevant indicators of abnormal cell division, yet their reliable detection remains challenging due to morphological ambiguity and scanner variability. In this work, we investigated three variants of adapting the pathology foundation model UNI2 for the MIDOG2025 Track 2 challenge: (1) LoRA + UNI2, (2) VPT + UNI2 + Vahadane Normalizer, and (3) VPT + UNI2 + GRL + Stain TTA. We observed that the integration of Visual Prompt Tuning (VPT) with stain normalization techniques contributed to improved generalization. The best robustness was achieved by further incorporating test-time augmentation (TTA) with Vahadane and Macenko stain normalization. Our final submission achieved a balanced accuracy of 0.8837 and an ROC-AUC of 0.9513 on the preliminary leaderboard, ranking within the top 10 teams. These results suggest that prompt-based adaptation combined with stain-normalization TTA offers a promising strategy for atypical mitosis classification under diverse imaging conditions.
@@ 0.3 Implementation Details
During inference, we adopt a test-time augmentation (TTA) strategy to improve the robustness of predictions. Each test patch is evaluated under multiple transformations, including horizontal and vertical flips as well as $90^{\circ}$ rotations. In addition to color augmentations, we apply two complementary stain normalization pipelines, namely Vahadane [10] and Macenko [11]. For each input, predictions across all augmented versions are averaged to obtain the final probability. This ensemble-like strategy reduces prediction variance and improves generalization across scanners.
@@ Results
VPT + UNI2-h + VahadaneNorm 0.8711 0.9014 0.8408 0.9483
VPT + UNI2-h + GRL + Stain TTA 0.8837 0.9577 0.8097 0.9513
We compared three variants of our framework on the MI-DOG2025 Track 2 dataset. As shown in Table 1, the LoRA + UNI2-h baseline achieved a balanced accuracy of 0.8305 and an ROC-AUC of 0.9364. Replacing LoRA with Visual Prompt Tuning (VPT) along with Vahadane stain normalization improved the results, yielding a balanced accuracy of 0.8711 and an ROC-AUC of 0.9483. Finally, incorporating a Gradient Reversal Layer (GRL) for domain adaptation and test-time augmentation (TTA) with both Vahadane and Macenko normalization further enhanced robustness, achieving a balanced accuracy of 0.8837 and the highest ROC-AUC of 0.9513. Based on these results, we selected the variant combining VPT, GRL, and TTA with UNI2-h as our final submission.
@@ Conclusion
In this study, we investigated three variants of adapting the UNI2-h foundation model for atypical mitosis classification in the MIDOG2025 Track 2 challenge. The LoRA-based approach or VPT provided a competitive baseline, while stain normalization improved generalization. Further integrating GRL along with test-time augmentation yielded the best results, achieving a balanced accuracy of 0.8837 and an ROC-AUC of 0.9513 on the preliminary test set. These findings demonstrate that combining prompt-based adaptation, stain normalization, and adversarial domain adaptation is an effective strategy for robust mitosis classification under diverse imaging conditions.